OpenStudy (anonymous):

Evaluate 3{5 + 3[10 + 4 · 8]} is it 393?

OpenStudy (anonymous):

this is about precedence: BODMAS(brackets,Of,division, multipication,add,subtract) 10+4=14-8=6 3(5+3(6)) 3(5+18) 3(23) 69.